The supplement of an angle is 10 degrees less than 3 times its compliment. Find the ratio of supplement to its compliment.
Answer by addingup(3677)

I don't know what you mean by compliment. Angles are either COMPLEMENTARY or SUPPLEMENTARY.
Supplementary angles are those that add up to 180 degrees.
Let's call one angle x, and the problem says the supplement is 3(x - 10):
x + 3(x - 10) = 180 Simplify
x + 3x - 30 = 180 Add x's on left. Also add 30, both sides.
4x = 210
x = 210/4
x = 52 1/2 degrees. And the other is 3(52.5-10)= 127 1/2 degrees